OpenStudy (anonymous):

How to factor this? 7x3 + 56 Thank you!

OpenStudy (samanthagreer):

7(x^3 - 8) There is a formula for this, it's called the difference of cubes a^3 - b^3 = ( a - b ) ( a^2 + ab + b^2 ) Is is obviously x and you should know that 2^3 = 8 so just plug that in 7(x - 2) (x^2 + 2x +4)

OpenStudy (anonymous):

Thank you so much, @samanthagreer !

OpenStudy (samanthagreer):

can i have a medal please np